OpenStudy (anonymous):

a clerk can process the mailing in 9 days, a new clerk takes 14 days to process them. if they work together, how many days will it take to do the processing?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

clerk A= 9 day so 1 day wrk=1/9 clerk B=14 day so 1 day wrk=1/14 together 1 day wrk =>1/9+1/14 =23/126 both can complete wrk in 126/3 day =

OpenStudy (anonymous):

sry 126/23 days

OpenStudy (anonymous):

thanks!!!!!

OpenStudy (anonymous):

correct ... medel plz :)

OpenStudy (anonymous):

how do i do that >< sorry i'm new